¿còmo derivo (x^2)x(valor absoluto de x^3)
Answer by stanbon(57307)

If X>=0 X^2*|X^3|= X^2*X^3 = X^5
If X<0 X^2*|X^3|= X^2*-X^3= -X^5
Cheers,
Stan H.